# Creativity and Machine Learning: A Survey

### v4

@@ 3.5.2. Examples of Models
Diffusion models have been primarily used for image generation. In order to generate higher-quality images and to allow text-to-image generation, a variety of effective methods for conditioning have been proposed. A possibility is to use classifier guidance (Dhariwal and Nichol, 2021): the diffusion score (i.e., the added noise) includes the gradient of the log-likelihood of an auxiliary classifier model. An alternative is classifier-free guidance (Ho and Salimans, 2021): to avoid learning an additional model, a single neural network is used to parameterize two diffusion models, one conditional and one unconditional; the two models are then jointly trained by randomly setting the class for the unconditional model. Finally, the sampling is performed using a linear combination of conditional and unconditional score estimates. Guided Language to Image Diffusion for Generation and Editing (GLIDE) (Nichol et al., 2022a) demonstrates how classifier-free guidance can be effectively used to generate text-conditional images. In addition, it shows how diffusion models can be used for image editing by fine-tuning in order to reconstruct masked regions. Performance improvement can be obtained by means of a cascade of multiple diffusion models performing conditioning augmentation (Ho et al., 2022a). Notably, the diffusion model can operate on latent vectors instead of real images. For example, Stable Diffusion (Rombach et al., 2022) employs a diffusion model in the latent space of a pre-trained autoencoder. Similarly, DALL-E 2 (Ramesh et al., 2022) generates images by conditioning with image representations. At first, it learns a prior diffusion model to generate possible CLIP image embeddings from a given text caption, i.e., conditioned by its CLIP text embedding. Then, a diffusion decoder produces images conditioned by the image embedding. The generation quality can be further improved by means of better, generated captions for the images in the training set (Betker et al., 2024). Imagen (Saharia et al., 2022) uses instead a cascaded diffusion decoder, together with a frozen language model as text encoder to increase the quality of output.
@@
Although the approach is particularly suitable to images, applications to other data sources have been developed as well. For example, DiffWave (Kong et al., 2021) and WaveGrad (Chen et al., 2021) use diffusion models to generate audio. They overcome the continuous-discrete dichotomy by working on waveform. Another possibility is to use an auto-encoder like MusicVAE (Roberts et al., 2018) to transform the sequence into a set of continuous latent vectors, on which training a diffusion model (Mittal et al., 2021). Resembling image generators, Contrastive Language-Audio Pretraining (CLAP) embeddings (Elizalde et al., 2022) can be used to generate audio by conditioning on text descriptions (Liu et al., 2023). Diffusion-LM (Li et al., 2022) employs diffusion models to write text by denoising a sequence of Gaussian vectors into continuous word vectors (then converted into discrete words by a rounding step); DiffuSeq (Gong et al., 2023) performs sequence-to-sequence generation tasks by embedding source and target sequences in the same embedding space through a Transformer architecture. Diffusion models have been used for 3D generation as well (Nichol et al., 2022b). Finally, diffusion models for video have also been proposed, based on gradient-based conditioning (Ho et al., 2022b), and on processing latent spacetime patches. In particular, with respect to the latter, Sora (Brooks et al., 2024) first turns videos into sequences of patches and then uses a diffusion Transformer to predict the original patches from random noise (and conditioning inputs like text prompts), improving sample quality and flexibility.
@@ 3.5.3. Applications
Despite their recent introduction, diffusion models have been used to generate audio, music, and video, as well as to generate and edit images conditioned on input text, e.g., with in-painting (Lugmayr et al., 2022) or subject-driven generation (Ruiz et al., 2022); we refer to (Yang et al., 2023) for a comprehensive survey of this area. Indeed, they lead to higher-quality outputs than the previous state-of-the-art models. In particular, DALL-E 2 and Stable Diffusion have been able to produce images from textual instructions with superior fidelity and variety.
@@ 3.6.1. Core Concepts
With RL-based methods we aim to indicate all the generative models whose training relies on maximizing a reward. These models are based on the architectures introduced so far, e.g., they can be GANs or autoregressive models. The difference is that they are not (only) trained to fool the discriminative part or to reduce prediction error. The typical framework considers the generative model as the agent; each action causes a modification to the current product, i.e., the state; and the agent learns a policy that maximizes the cumulative reward. Therefore, the reward can be used to impose desired behavior to the generative model. The RL-based approach can be implemented for the entire training or for fine-tuning a pre-trained model. The sampling scheme remains the same, and depends on the chosen generative model.
@@ 3.6.2. Examples of Models
We have already introduced a fully RL-trained model in Section 3.2: ORGAN (Guimaraes et al., 2017). Here, RL is used to both adapt GANs to sequential tasks, and to provide additional learning signals, such as rewards from specific-domain objectives. Also (Yi et al., 2018) follows this path by using rewards like fluency, coherence, meaningfulness, and overall quality to generate poems. Another possibility is to use the metrics used at test time (e.g., BLEU or ROUGE) (Bahdanau et al., 2017; Ranzato et al., 2016). Instead, RL-DUET (Jiang et al., 2020) casts online music accompaniment generation as a RL problem with an ensemble of reward models, i.e., autoregressive models trained with or without the whole context, and with or without the human-produced counterpart. In this way, inter-coherence (between humans and machines) and intra-coherence can be obtained. Finally, Intelli-Paint (Singh et al., 2022) can paint in human style by using a sequential planner that learns a painting policy to predict vectorized brushstroke parameters from the current state.
@@
RL can also be used to fine-tune a pre-trained generative model. Doodle-SDQ (Zhou et al., 2018) first learns to draw simple strokes using supervised learning; then, it improves its generation by means of rewards about similarity, color, and line movement. Conversely, the authors of (Tambwekar et al., 2019) suggest to consider a pre-trained LSTM language model as a policy model. Fine-tuning then aims at maximizing the probability that a given event occurs at the end of the narrative. RL Tuner (Jaques et al., 2017b) uses RL to fine-tune a Note-RNN (Eck and Schmidhuber, 2002) to produce music that follows a set of music theory rules. To avoid forgetting note probabilities learned from the data, the probability value returned by a copy of the pre-trained Note-RNN can be used as an additional reward. Sequence Tutor (Jaques et al., 2017a) generalizes this idea of learning a policy that trades off staying close to the data distribution while improving performance on specific metrics. A comprehensive critical discussion of the rewards for RL-based generative models can be found in (Franceschelli and Musolesi, 2024). Finally, RL can be used to help models follow human preferences (Christiano et al., 2017) or feedback. The latter technique is referred to as Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020). For example, ChatGPT [^5] https://openai.com/blog/chatgpt/ is an interactive version of GPT-3 (Brown et al., 2020) (initially) and GPT-4 (OpenAI, 2023) (at the time of writing), fine-tuned to maximize a learned reward of human values. RLHF improves its conversational skills while mitigating mistakes and biases; because of this, it has become a standard de facto for fine-tuning large language models, e.g., in (Touvron et al., 2023b). It is also possible to use AI and not human feedback (Bai et al., 2022; Lee et al., 2023). However, RLHF has some limitations (Casper et al., 2023), and alternative RL-free strategies are increasingly popular (e.g., (Rafailov et al., 2023)).
@@ 3.6.4. Critical Discussion
The evaluation of creativity of RL-based models depends on how the agent is implemented and which rewards are considered. The learned space of solutions depends on the used rewards (and on the pre-training technique in case of fine-tuning). They typically contain an adversarial signal or a likelihood with respect to the training data; thus, combinatorial or exploratory creativity is obtained. However, additional rewards can have the effect of transforming that space (see Section 3.8). As far as Boden’s criteria are concerned, value is typically ensured by some qualitative domain-specific reward or by approaching human preferences. This also allows us to consider them as appreciative, as discussed in Section 3.2.4. Novelty and surprise might be achieved as well by means of specific rewards; however, this is not the case for current models.
@@ 3.7.1. Core Concepts
The last class of methods we consider in our analysis is not about a different generative model. On the contrary, it is about a different way to sample results from (pre-trained) generative models, namely by means of its inputs. Two different approaches can be used. The first is about carefully selecting or optimizing the input to a generative model (e.g., the latent vector or the text prompt), so that to obtain the desired output. The second approach is about optimizing the input, so that it directly becomes the desired output. They rely on losses that are usually based on features learned by neural networks. While the two approaches are technically different, both of them aim at obtaining better outputs by exploiting the knowledge of the pre-trained model through the optimization of the inputs.
@@ 3.7.2. Examples of Models
The first approach consists in carefully modifying the input of a generative model until the output matches the desired properties. The main example is VQGAN-CLIP (Crowson et al., 2022). Given a text description, VQGAN produces a candidate image from a random latent vector; the vector is then optimized by minimizing the distance between the embeddings of the description and the candidate image. Both embeddings are computed using CLIP (Radford et al., 2021). Variants can be implemented as in Wav2CLIP (Wu et al., 2022), where an audio encoder is learned to match the CLIP encoders so that VQGAN-CLIP can be used from raw audio; or as in music2video (Jang et al., 2022), where videos are generated from audio a frame after the other by both minimizing the distance between subsequent frames, and the distance between image and music segment embedded by Wav2CLIP. In addition to the random latent vector, the text or audio description can be optimized as well. This can be performed by the users through many iterations of careful adjustments, or by means of an automated procedure. The latter is commonly known as prompt tuning. Prompt tuning is about producing prompts via backpropagation; the optimized prompts can then condition frozen language models in order to perform specific tasks without having to fine-tune them (Lester et al., 2021). An additional model can also be trained to output the desired prompt (Levine et al., 2022). Finally, image generators such as VQGAN can also be exploited in other ways, i.e., with binary-tournament genetic algorithm (Fernando et al., 2021) or more complex evolution strategies (Tian and Ha, 2022). Another possibility is to optimize the input so that the generated output maximizes a target neuron of an image classifier (Nguyen et al., 2016). This helps generate what that neuron has learned. The desired latent vector can also be produced by an additional model (Nguyen et al., 2017).
The second approach consists in optimizing the inputs so that to transform them into the desired outputs. DeepDream (Mordvintsev et al., 2015) generates “hallucinated” images by modifying the input to maximize the activation of a certain layer from a pre-trained classifier. Artistic style transfer is based on the same idea. Given an input image and a target image, the former is modified by means of both style and content losses thanks to a pre-trained classifier. The content loss is minimized if the current and the original input images generate the same outputs from the hidden layers. The style loss is minimized if the current and target images have the same pattern of correlation between feature maps in the hidden layers (Gatys et al., 2016). Control over results can be improved by considering additional losses about color, space, and scale (Gatys et al., 2017).
@@ 3.7.3. Applications
Input-based methods can be used with any generative model to produce the desired output. With language models, they can exploit their generality in several specific tasks without fine-tuning them. For instance, prompt tuning can be used by writers for co-creation (Chakrabarty et al., 2023) or to force LLMs to brainstorm (Summers-Stay et al., 2023). With image generators, they can obtain drawings adherent to given descriptions, or high-quality but yet peculiar paintings like colorist (Fernando et al., 2021), abstract (Tian and Ha, 2022) or alien (Snell, 2021) artworks. Videos can be generated too [^6] Also artistic abstract videos can be produced by navigating the latent space: www.jakeelwes.com/project-latentSpace.html, and we believe applications to other domains are yet to come. Both types of input-based methods can be used not only to produce desired outputs or to transfer styles; they can also be used to better analyze what it is inside the network (Nguyen et al., 2016; Olah et al., 2017).
@@ 3.7.4. Critical Discussion
Since input-based methods are applied on pre-trained generative models, the space of solutions in which they work is the one induced by those models, i.e., the common spaces we can derive from real data. Nonetheless, some techniques may be able to cause productions that are outside that space or at its boundaries, i.e., to cause transformational creativity. This can especially happen if the model is general, and the output for a specific task is not only sampled from the sub-space of solutions for that task (e.g., with prompt tuning over a language model).
Input-based methods are also valuable: the input optimization itself is typically guided by some sort of qualitative loss. On the other hand, they are not explicitly novel or surprising (although the results might seem so). However, nothing prevents to optimize the loss in such directions (see Section 3.8).
@@ 3.8. Practical Assessment of Creativity-Oriented Methods
Another possibility is the development of an input-based method where the input is optimized so that to obtain a product that is valuable, novel, and surprising. This may be achieved either by forcing a further exploration of the latent space (e.g., by means of evolutionary search (Fernandes et al., 2020)), or by defining appropriate loss functions to perform gradient descent over the input. All these methodologies are also called active divergence (Berns and Colton, 2020), since they aim to generate in ways that do not simply reproduce training data. A survey on active divergence can be found in (Broad et al., 2021). A different output can also be obtained by carefully altering the probability distribution of the model, e.g., by scaling its probabilities with learned functions to maximize target properties (Dathathri et al., 2020; Snell et al., 2023; Yang and Klein, 2021).
A different approach is followed by the Composer-Audience architecture (Bunescu and Uduehi, 2019). Two models are considered: the Audience, a simple sequence prediction model trained on a given dataset; the Composer, another sequence prediction model trained on a different dataset. In addition, the Composer also receives the next-token expectations from the Audience, and it learns when to follow its guidance and when to diverge from expectations, i.e., when to be surprising. For instance, it can learn to produce jokes by considering non-humorous texts to train the Audience, and humorous texts to train the Composer. Even though this approach is useful to learn how to generate valuable and surprising output, it is only applicable when paired datasets are available.

### v6

@@ 3.5.2. Examples of Models
Diffusion models have been primarily used for image generation. In order to generate higher-quality images and to allow text-to-image generation, a variety of effective methods for conditioning have been proposed. A possibility is to use classifier guidance (Dhariwal and Nichol, 2021): the diffusion score (i.e., the added noise) includes the gradient of the log-likelihood of an auxiliary classifier model. An alternative is classifier-free guidance (Ho and Salimans, 2021): to avoid learning an additional model, a single neural network is used to parameterize two diffusion models, one conditional and one unconditional; the two models are then jointly trained by randomly setting the class for the unconditional model. Finally, the sampling is performed using a linear combination of conditional and unconditional score estimates. Guided Language to Image Diffusion for Generation and Editing (GLIDE) (Nichol et al., 2022a) demonstrates how classifier-free guidance can be effectively used to generate text-conditional images. In addition, it shows how diffusion models can be used for image editing by fine-tuning in order to reconstruct masked regions. Performance improvement can be obtained by means of a cascade of multiple diffusion models performing conditioning augmentation (Ho et al., 2022a). Notably, the diffusion model can operate on latent vectors instead of real images. Stable Diffusion (Rombach et al., 2022) employs a diffusion model in the latent space of a pre-trained autoencoder. Similarly, DALL-E 2 (Ramesh et al., 2022) generates images by conditioning with image representations. At first, it learns a prior diffusion model to generate possible CLIP image embeddings from a given text caption, i.e., conditioned by its CLIP text embedding. Then, a diffusion decoder produces images conditioned by the image embedding. The generation quality can be further improved by means of generated captions for the images in the training set (Betker et al., 2024). Imagen (Saharia et al., 2022) uses instead a cascaded diffusion decoder, together with a frozen language model as a text encoder to increase the quality of output.
@@
Although the approach is particularly suitable for images, applications to other data sources have been developed as well. DiffWave (Kong et al., 2021) and WaveGrad (Chen et al., 2021) use diffusion models to generate audio. They overcome the continuous-discrete dichotomy by working on waveform. Another possibility is to use an auto-encoder like MusicVAE (Roberts et al., 2018) to transform the sequence into a set of continuous latent vectors, on which training a diffusion model (Mittal et al., 2021). Resembling image generators, Contrastive Language-Audio Pretraining (CLAP) embeddings (Elizalde et al., 2023) can be used to generate audio by conditioning on text descriptions (Liu et al., 2023). Diffusion-LM (Li et al., 2022) employs diffusion models to write text by denoising a sequence of Gaussian vectors into continuous word vectors (then converted into discrete words by a rounding step); DiffuSeq (Gong et al., 2023) performs sequence-to-sequence generation tasks by embedding source and target sequences in the same embedding space through a Transformer architecture. Diffusion models have been used for 3D generation as well (Nichol et al., 2022b). Finally, diffusion models for video have also been proposed, based on gradient-based conditioning (Ho et al., 2022b), and on processing latent spacetime patches. In particular, with respect to the latter, Sora (Brooks et al., 2024) first turns videos into sequences of patches and then uses a diffusion Transformer to predict the original patches from random noise (and conditioning inputs like text prompts), improving sample quality and flexibility.
@@ 3.5.3. Applications
Despite their recent introduction, diffusion models have been used to generate audio, music, and video, as well as to generate and edit images conditioned on input text, e.g., with in-painting (Lugmayr et al., 2022) or subject-driven generation (Ruiz et al., 2023); we refer to (Yang et al., 2023) for a comprehensive survey of this area. Indeed, they lead to higher-quality outputs than the previous state-of-the-art models. In particular, DALL-E 2 and Stable Diffusion have been able to produce images from textual instructions with superior fidelity and variety.
@@ 3.6.1. Core Concepts
With RL-based methods, we aim to indicate all the generative models whose training relies on maximizing a reward. These models are based on the architectures introduced so far, e.g., they can be GANs or autoregressive models. The difference is that they are not (only) trained to fool the discriminative part or to reduce prediction error. The typical framework considers the generative model as the agent; each action causes a modification to the current product, i.e., the state; and the agent learns a policy that maximizes the cumulative reward. Therefore, the reward can be used to impose desired behavior on the generative model. The RL-based approach can be implemented for the entire training or for fine-tuning a pre-trained model. The sampling scheme remains the same and depends on the chosen generative model.
@@ 3.6.2. Examples of Models
A first example is Objective-Reinforced GAN (ORGAN) (Guimaraes et al., 2017). Here, RL is used not only to adapt GANs to sequential tasks but also to provide additional learning signals, such as rewards from specific-domain objectives (e.g., tonality and ratio of steps for music generation). Also (Yi et al., 2018) follows this path by using rewards like fluency, coherence, meaningfulness, and overall quality to generate poems. Another possibility is to use the metrics used at test time (e.g., BLEU or ROUGE) (Bahdanau et al., 2017; Ranzato et al., 2016). Instead, RL-DUET (Jiang et al., 2020) casts online music accompaniment generation as an RL problem with an ensemble of reward models, i.e., autoregressive models trained with or without the whole context, and with or without the human-produced counterpart. In this way, inter-coherence (between humans and machines) and intra-coherence can be obtained. Finally, Intelli-Paint (Singh et al., 2022) can paint in human style by using a sequential planner that learns a painting policy to predict vectorized brushstroke parameters from the current state.
@@
RL can also be used to fine-tune a pre-trained generative model. Doodle-SDQ (Zhou et al., 2018) first learns to draw simple strokes using supervised learning; then, it improves its generation by means of rewards about similarity, color, and line movement. Conversely, the authors of (Tambwekar et al., 2019) suggest to consider a pre-trained LSTM language model as a policy model. Fine-tuning then aims at maximizing the probability that a given event occurs at the end of the narrative. RL Tuner (Jaques et al., 2017b) uses RL to fine-tune a Note-RNN (Eck and Schmidhuber, 2002) to produce music that follows a set of music theory rules. To avoid forgetting note probabilities learned from the data, the probability value returned by a copy of the pre-trained Note-RNN can be used as an additional reward. Sequence Tutor (Jaques et al., 2017a) generalizes this idea of learning a policy that trades off staying close to the data distribution while improving performance on specific metrics. A comprehensive critical discussion of the rewards for RL-based generative models can be found in (Franceschelli and Musolesi, 2024). Finally, RL can be used to help models follow human preferences (Christiano et al., 2017) or feedback. The latter technique is referred to as Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020). For example, ChatGPT (OpenAI, 2022) is an interactive version of GPT-3 (Brown et al., 2020) (initially) and GPT-4 (OpenAI, 2023) (at the time of writing), fine-tuned to maximize a learned reward of human values. RLHF improves its conversational skills while mitigating mistakes and biases; because of this, it has become a standard de facto for fine-tuning large language models, e.g., in (Touvron et al., 2023b). It is also possible to use AI and not human feedback (Bai et al., 2022; Lee et al., 2023). However, RLHF has some limitations (Casper et al., 2023), and alternative RL-free strategies are increasingly popular (e.g., (Rafailov et al., 2023)).
@@ 3.6.4. Critical Discussion
The evaluation of the creativity of RL-based models depends on how the agent is implemented and which rewards are considered. The learned space of solutions depends on the used rewards (and on the pre-training technique in case of fine-tuning). They typically contain an adversarial signal or a likelihood with respect to the training data; thus, combinatorial or exploratory creativity is obtained. However, additional rewards can have the effect of transforming that space (see Section 3.8). As far as Boden’s criteria are concerned, value is typically ensured by some qualitative domain-specific reward or by approaching human preferences. This also allows us to consider them as appreciative, as discussed in Section 3.2.4. Novelty and surprise might be achieved as well by means of specific rewards; however, this is not the case for current models.
@@ 3.7.1. Core Concepts
The last class of methods we consider in our analysis is not about a different generative model. On the contrary, it is about a different way to sample results from (pre-trained) generative models, namely by means of its inputs. Two different approaches can be used. The first is about carefully selecting or optimizing the input to a generative model (e.g., the latent vector or the text prompt) so that to obtain the desired output. The second approach is about optimizing the input so that it directly becomes the desired output. They rely on losses that are usually based on features learned by neural networks. While the two approaches are technically different, both of them aim at obtaining better outputs by exploiting the knowledge of the pre-trained model through the optimization of the inputs.
@@ 3.7.2. Examples of Models
The first approach consists of carefully modifying the input of a generative model until the output matches the desired properties. The main example is VQGAN-CLIP (Crowson et al., 2022). Given a text description, VQGAN produces a candidate image from a random latent vector; the vector is then optimized by minimizing the distance between the embeddings of the description and the candidate image. Both embeddings are computed using CLIP (Radford et al., 2021). Variants can be implemented as in Wav2CLIP (Wu et al., 2022), where an audio encoder is learned to match the CLIP encoders so that VQGAN-CLIP can be used from raw audio; or as in music2video (Jang et al., 2022), where videos are generated from audio a frame after the other by both minimizing the distance between subsequent frames, and the distance between image and music segment embedded by Wav2CLIP. In addition to the random latent vector, the text or audio description can be optimized as well. This can be performed by the users through many iterations of careful adjustments, or by means of an automated procedure. The latter is commonly known as prompt tuning. Prompt tuning is about producing prompts via backpropagation; the optimized prompts can then condition frozen language models in order to perform specific tasks without having to fine-tune them (Lester et al., 2021). An additional model can also be trained to output the desired prompt (Levine et al., 2022). Finally, image generators such as VQGAN can also be exploited in other ways, i.e., with binary-tournament genetic algorithm (Fernando et al., 2021) or more complex evolution strategies (Tian and Ha, 2022). Another possibility is to optimize the input so that the generated output maximizes a target neuron of an image classifier (Nguyen et al., 2016). This helps generate what that neuron has learned. The desired latent vector can also be produced by an additional model (Nguyen et al., 2017).
The second approach consists of optimizing the inputs to transform them into the desired outputs. DeepDream (Mordvintsev et al., 2015) generates “hallucinated” images by modifying the input to maximize the activation of a certain layer from a pre-trained classifier. Artistic style transfer is based on the same idea. Given an input image and a target image, the former is modified by means of both style and content losses thanks to a pre-trained classifier. The content loss is minimized if the current and the original input images generate the same outputs from the hidden layers. The style loss is minimized if the current and target images have the same pattern of correlation between feature maps in the hidden layers (Gatys et al., 2016). Control over results can be improved by considering additional losses about color, space, and scale (Gatys et al., 2017).
@@ 3.7.3. Applications
Input-based methods can be used with any generative model to produce the desired output. With language models, they can exploit their generality in several specific tasks without fine-tuning them. For instance, prompt tuning can be used by writers for co-creation (Chakrabarty et al., 2023) or to force LLMs to brainstorm (Summers-Stay et al., 2023). With image generators, they can obtain drawings adherent to given descriptions, or high-quality but yet peculiar paintings like colorist (Fernando et al., 2021), abstract (Tian and Ha, 2022) or alien (Snell, 2021) artworks. We believe applications to other domains are yet to come. Both types of input-based methods can be used not only to produce desired outputs or to transfer styles; they can also be used to better analyze what is inside the network (Nguyen et al., 2016; Olah et al., 2017).
@@ 3.7.4. Critical Discussion
Since input-based methods are applied to pre-trained generative models, the space of solutions in which they work is the one induced by those models, i.e., the common spaces we can derive from real data. Nonetheless, some techniques may be able to cause productions that are outside that space or at its boundaries, i.e., to cause transformational creativity. This might happen if the model is general, and the output for a specific task is not only sampled from the sub-space of solutions for that task (e.g., with prompt tuning over a language model). Input-based methods are also valuable: the input optimization itself is typically guided by some sort of qualitative loss. On the other hand, they are not explicitly novel or surprising (although the results might seem so). However, nothing prevents optimizing the loss in such directions (see Section 3.8).
@@ 3.8. Practical Assessment of Creativity-Oriented Methods
Another possibility is the development of an input-based method where the input is optimized to obtain a product that is valuable, novel, and surprising. This may be achieved either by forcing a further exploration of the latent space (e.g., by means of evolutionary search (Fernandes et al., 2020)), or by defining appropriate loss functions to perform gradient descent over the input. All these methodologies are also called active divergence (Berns and Colton, 2020) since they aim to generate in ways that do not simply reproduce training data. A survey on active divergence can be found in (Broad et al., 2021). A different output can also be obtained by carefully altering the probability distribution of the model, e.g., by scaling its probabilities with learned functions to maximize target properties (Dathathri et al., 2020; Snell et al., 2023; Yang and Klein, 2021).
A different approach is followed by the Composer-Audience architecture (Bunescu and Uduehi, 2019). Two models are considered: the Audience, a simple sequence prediction model trained on a given dataset; the Composer, another sequence prediction model trained on a different dataset. In addition, the Composer also receives the next-token expectations from the Audience, and it learns when to follow its guidance and when to diverge from expectations, i.e., when to be surprising. For instance, it can learn to produce jokes by considering non-humorous texts to train the Audience, and humorous texts to train the Composer. Even though this approach is useful for learning how to generate valuable and surprising output, it is only applicable when paired datasets are available.